how do i solve (2t-3)^2-4(2t-3)+3=0
Answer by scott8148(6628)

factoring ___ [(2t - 3) - 3][(2t - 3) - 1] = 0

(2t - 3) - 3 = 0

(2t - 3) - 1 = 0